at opposee comen oatquare whose sides have a length of 0.330 m. A test charge (ao -2.30 x 10a C), with a mass of 1.10 x 10-8 s released from rest at one of the corners of the square. Determine the speed of the test charge when it reaches the center of the square Number Units
Answer #1

V1 = 2k q/r = 2 x 9 x 10^9 x 7.5 x 10^-6/0.33 = 409090.91 V

V2 = 4 x 9 x 10^9 x 7.5 x 10^-6/(0.33 x sqrt(2)) = 578541.91 V

dV = V2 - V1 = 1.69 x 10^5 V

Work, W = q dV = 0.5 mv^2

Velocity, v = sqrt(2q dV/m)

V = sqrt(2 x 2.3 x 10^-8 x 1.69 x 10^5/1.1 x 10^-8)

V = 841.79 m/s